OpenStudy (anonymous):

solve for the derivative og the folowing function implicitly x^3 - xy + y^2 =7 help me to solve this guys.

OpenStudy (anonymous):

@ayo123 help

OpenStudy (anonymous):

ok, where dy/dx =y' D(x^3 - xy + y^2) =D(7) 3x^2 - ((1)y + y' x) + 2y y' = 0 3x^2 - (y + y' x) + 2y y' = 0 3x^2 - y - y'x + 2yy' = 0 collecting like terms, - y'x + 2yy'= y- 3x^2 y' (2y-x) = y- 3x^2 y' =(y- 3x^2) /(2y-x)

OpenStudy (anonymous):

@ayo123its that it?

OpenStudy (anonymous):

@ayo123 its that it|?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@iGreen help

OpenStudy (anonymous):

@hartnn help

OpenStudy (anonymous):

dy/dx = (3y - 2x) / (2y - 3x)